OpenStudy (anonymous):

The area of a rectangle is given by A = x2 + 4x – 5. Find the value of x if the total area of the rectangle is equal to 7 square units. a) x = 1 b) x = 7 c) x = 2 d) x = 5

OpenStudy (anonymous):

c

OpenStudy (radar):

\[7=x ^{2}+4x-5\] \[x ^{2}+4x-12=0\]factoring would be the way to go! (x+6)(x-2)=0 If this product equals zero then one (or both) equals zero x+6=0 giving x=-6 not a good answer as dimensions are positive x-2=0 giving x=2 That is a good solution. But better check and verify! \[7=2^{2}+4(2)-5\]\[7=4+8-5=7\] Final answer x=2